OpenStudy (calculusxy):

The sum of the first 5 terms of an arithmetic series is 170. The sum of the first 6 terms is 225. The common difference is 7. Determine the first 4 terms of the series.

OpenStudy (calculusxy):

@ganeshie8 @jim_thompson5910 @satellite73

rishavraj (rishavraj):

\[S_5 = 2a_1 + (5 - 1)d\] \[S-6 = 2a_1 + (6 - 1)d\] u got d .....and S_5 &S_6 so u can easily get a_1 i.e the first term and after tht get the first four terms

OpenStudy (calculusxy):

I have learned the summation formula to be \(S_n=\frac{n}{2}(a_1+a_2}. Is this different from what @rishavraj posted?

jimthompson5910 (jim_thompson5910):

S5 = sum of the first five terms S6 = sum of the first six terms In this case, we're told S5 = 170 S6 = 225 Subtracting the sums gives S6 - S5 = 225 - 170 = 55 which is the 6th term. This is because S6 = (sum of first five terms) + (sixth term)

rishavraj (rishavraj):

@calculusxy oops no I forgot to mention n/2

OpenStudy (calculusxy):

\(S_n = \frac{n}{2}(a_1 + a_n)\)

OpenStudy (calculusxy):

So then we would have a_1 + 7(5) = 55 a_1 = 20

OpenStudy (calculusxy):

And then 20, 27, 34, 41,...

rishavraj (rishavraj):

yup

OpenStudy (calculusxy):

Thanks. And I have noticed that in some problems of arithmetic series, there is a 2a_1 and d(n-1). This is different from what I have learned and I would like to know the difference between them if there is any.

rishavraj (rishavraj):

see u know \[S_n = \frac{ n }{ 2 }(a_1 + a_n)\] \[a_n = a_1 + (n-1)d\] so we plug tht value we get \[S_n = \frac{ n }{ 2 }(a_1 + a_1 + (n - 1)d)\]

OpenStudy (calculusxy):

So basically the second part of the terms in the parenthesis where there's a_1 repeated and so on, it just is the formula for finding the a_n? Thus, it is the same thing? And when would I use this formula that has the d with it as opposed to the other one?

rishavraj (rishavraj):

see if u got we say S_5 and S_8 bt not the common difference 'd' or the first term 'a' we only kgot the value of S_5 and S_8 we use tht formula to find a_1 and d ..because in tht case we would be having 2 variables i.e a_1 and d and 2 equations :)

OpenStudy (calculusxy):

Ok got it! Thanks :)

rishavraj (rishavraj):

yw :)))